# Focus On This, Not That! Steering LLMs With Adaptive Feature Specification

### v2

@@ 1 Introduction
Instruction Tuning (IT) (Zhang et al., 2023), a specialised form of supervised fine-tuning (SFT), has become an essential step in the process of developing effective instruction-following large language models (LLMs) (Ouyang et al., 2022; Touvron et al., 2023; Chen et al., 2024). While extensive pre-training to perform next token prediction allows LLMs to extract common patterns and knowledge from large text corpora, IT fine-tunes these models on input-output pairs complemented by natural-language task instructions, teaching them to perform open-ended language-based tasks given instructions (Huang et al., 2023).
@@
However, despite the improvements observed in zero-shot generalisation from IT, recent studies suggest that some of these gains may be superficial, stemming from the models’ ability to learn task template formats or spurious input/output correlations rather than a more generalisable instruction-following capability (Kung & Peng, 2023; Ghosh et al., 2024). As a result, LLMs may fail to generalise to new contexts where the same templates or spurious correlations are not present (Kung & Peng, 2023). Furthermore, fine-tuning can inadvertently lead to safety misalignment, where models lose alignment with desired objectives and become more prone to generating harmful or undesirable outputs (Qi et al., 2023). This motivates the need for methods that produce adaptive models at inference time, enabling users to steer model behaviour based on evolving requirements or safety considerations.
@@
To address these limitations, we propose Focus Instruction Tuning (FIT), an extension of traditional IT where LLMs are fine-tuned with respect to an instruction specifying which features to focus on” or not focus on.” By training LLMs to condition responses on these focus specifications—enabling responses to the same input to differ based on the specified features—FIT allows end users to dynamically steer model behavior. This capability not only provides precise control over which features play a role or are ignored in performing a task but also adds an important layer of explainability. For instance, in Figure 1, we demonstrate how FIT can steer a model to ignore gender stereotypes and focus on task-relevant information, enabling it to correctly solve a question-answering task.
@@
In our experiments, we demonstrate that FIT enables precise control over model behavior by steering it to focus on task-relevant features while ignoring features specified as irrelevant or spurious (see Section 4). This flexibility allows FIT to address challenges such as mitigating the influence of demographic features in question-answering tasks, while also proving effective across diverse NLP tasks like sentiment analysis and natural language inference. Furthermore, we show that FIT is robust to distribution shifts over feature values and can generalize to new, unseen features, demonstrating its adaptability to dynamic user requirements and varying task contexts.
In sum, our primary contributions are as follows:
@@
    We introduce Focus Instruction Tuning (FIT), a method that allows users to flexibly and dynamically specify what features a model should or should not focus on when performing a task at inference time. FIT enables practitioners to incorporate domain-specific knowledge about core, spurious, or bias-relevant features in order to steer models according to the desired feature specification, whilst adding a level of explainability.
@@
    We show that FIT generalises with respect to (a) distribution shift over feature values and (b) focusing on new features not seen during training, further demonstrating Focus Instruction Tuning ’s adaptability.
Figure 1: Focus Instruction Tuning (FIT). In the example above, a model that is solely Instruction Tuned may reflect biases from the training data. For instance, in a question from BBQ (Parrish et al., 2022), when asked who posed a technical question at an engineering candidate’s interview involving both a man and a woman, the model might incorrectly answer ”the man” due to biases, despite evidence to the contrary. In contrast, a FIT model can ignore the gender feature and focus on the interview content, demonstrating steerability and adaptability at inference time.
@@ 2.1 Spurious Feature Learning
Deep neural networks, such as foundation models like LLMs, are susceptible to relying on spurious features present in the training dataset – i.e., input features that are correlated with outputs in the training distribution, but are not correlated in all test distributions (Ye et al., 2024). Relying on spurious features leads models to fail to generalise under distribution shifts where such correlations may no longer hold Wang et al. (2023a). Spurious features have been extensively studied in computer vision, encompassing features such as background colour (Xiao et al., 2021; Venkataramani et al., 2024; Arjovsky et al., 2019) or texture (Geirhos et al., 2018; Baker et al., 2018), and are also prevalent in many widely used NLP benchmarks (Sun et al., 2024; Borkan et al., 2019). For instance, the token SPIELBERG is spuriously correlated with positive sentiment in datasets like SST-2 (Socher et al., 2013b), meaning that models trained on SST-2 may learn to predict sentiment by leveraging these spurious features instead of more general sentiment features (Wang & Culotta, 2020). This reliance on non-causal features undermines the robustness of models in generalising to distribution shift.
A variety of approaches have been explored to detect and mitigate the effects of spurious feature learning, particularly under distribution shifts. Traditional approaches include prompt engineering (Sun et al., 2024), regularisation techniques (Arjovsky et al., 2019; Chew et al., 2024), and directly incorporating causal inference strategies (Wang & Culotta, 2020; 2021; Udomcharoenchaikit et al., 2022). Substantial work in mechanistic interpretability has also aimed to discover models’ representation and use of task-causal or spurious features: for instance, causal probing (which trains probing classifiers to recognise and modify supervised feature representations encoded by foundation models; see Belinkov, 2022; Canby et al., 2024; Davies & Khakzar, 2024) has been used to study how models leverage causal versus spurious features features in the context of a given task (Ravfogel et al., 2021; Lasri et al., 2022; Davies et al., 2023). Other works have leveraged unsupervised mechanistic interpretability methods, such as circuit discovery techniques (Wang et al., 2023b; Conmy et al., 2023) and sparse auto-encoders (Subramanian et al., 2018; Yun et al., 2021), to improve generalisation by discovering spurious features leveraged by models in performing a given task and ablating their use of these features (Gandelsman et al., 2024; Marks et al., 2024). Finally, concept removal methods locate and manipulate supervised feature representations corresponding to bias features encoded by foundation models in order to remove these features (Ravfogel et al., 2020; 2022; 2023; Iskander et al., 2023; Belrose et al., 2024; Kuzmin et al., 2024).
@@ 2.2 Controlling LLMs
Steering LLMs. While alignment techniques like Reinforcement Learning with Human Feedback (RLHF) (Bai et al., 2022) have helped create more aligned and less harmful models, these models can still be prompted to generate harmful responses (Carlini et al., 2024). Furthermore, when these models are fine-tuned—currently the predominant paradigm for practical applications of LLMs—they have been shown to lose alignment with safety objectives and become more prone to producing harmful outputs (Qi et al., 2023).
This, combined with the practical infeasibility of fully retraining large foundation models, has driven the need for inference-time steering of LLMs to address safety misalignment. Recent work has proposed the use of steering vectors to modify attention weights (Bhattacharjee et al., 2024; Panickssery et al., ; Li et al., 2024) and the application of sparse autoencoders (SAEs) (Cunningham et al., 2023) to control the internal knowledge selection behaviour of LLMs (Zhao et al., 2024), enabling safer and more adaptive test-time responses. However, these methods typically assume white-box access to model representations, which can be complex and challenging to work with in practice. Furthermore, they do not train models to internalise test-time adaptability as a feature. Instead, they rely on external interventions for steering, lacking mechanisms for models to directly adapt their behaviour in response to user instructions provided in natural language.
Our work addresses these limitations by explicitly training models to exhibit controllable and robust test-time adaptability, enabling users to dynamically steer model behaviour through simple, natural language instructions.
@@ 3 Methodology
Preliminaries. We consider a pre-trained, decoder-only large language model (LLM) $p_{\theta}$ that models the probability distribution over its vocabulary $\mathcal{V}$ autoregressively. For an input sequence $x=[x_{1},\dots,x_{L}]\in\mathcal{V}^{L}$, the joint probability of $x$ is given by $p_{\theta}(x)=\prod_{i=1}^{L}p_{\theta}(x_{i}|x_{<i})$, with $p_{\theta}(x_{1}|\emptyset)=p_{\theta}(x_{1})$. In traditional supervised learning, for a sample $(x,y)\sim\mathcal{D}$, the conditional likelihood of the output $y$ given input $x$ is $p_{\theta}(y|x)=\prod_{i=1}^{L}p_{\theta}(y_{i}|x,y_{<i})$, with $p_{\theta}(y_{1}|\emptyset)=p_{\theta}(y_{1})$; and in supervised fine-tuning (SFT) of LLMs, this manifests as minimising the negative log-likelihood (NLL) of $y$ given $x$.
In instruction tuning (IT) (Zhang et al., 2023), a form of SFT, an additional task instruction $I$ accompanies the input-output pair $(x,y)$, forming a tuple $(I,x,y)$. The objective becomes minimising the NLL of $y$ given both $I$ and $x$, i.e., $p_{\theta}(y|I,x)$.
@@
$\mathcal{I}_{\text{focus}}=\{{\color[rgb]{0,0.4453125,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4453125,0.69921875}\emptyset},\;{\color[rgb]{0.90234375,0.625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.625,0}\text{focus}(F_{i})},\;{\color[rgb]{0.80078125,0.47265625,0.65625}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0.47265625,0.65625}\text{ignore}(F_{j})},\;{\color[rgb]{0,0.62109375,0.44921875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.62109375,0.44921875}\text{focus}(F_{i})\wedge\text{ignore}(F_{j})}\mid F_{i},F_{j}\in\mathcal{F}\},$ (1)
where: $\emptyset$ denotes an empty focus instruction with no features to focus on or to ignore; $\text{focus}(F_{i})$ is an instruction to focus on feature $F_{i}$; $\text{ignore}(F_{j})$ is an instruction to ignore feature $F_{j}$; and $\text{focus}(F_{i})\wedge\text{ignore}(F_{j})$ is an instruction to focus on feature $F_{i}$ whilst ignoring feature $F_{j}$. We include the default prompt in order to aid the model in learning the underlying task as well as the ability to refocus its attention on specified features during FIT. For specific examples of focus instrucitons that we consider, see Section A.4.
Consider a sample $(x,y)\sim p_{\text{data}}(x,y)$ drawn from an underlying data distribution and a focus instruction $I_{\text{focus}}$ drawn from a distribution $p_{\mathcal{I}_{\text{focus}}}$ over the set of focus instructions $\mathcal{I}_{\text{focus}}$. Then the likelihood of response $y$ conditioned on input $x$, task instruction $I$ (as in standard IT), and focus-instruction $I_{\text{focus}}$ is modelled as $p_{\theta}(y|I,I_{\text{focus}},x)$.
@@
For a sample $(x,y)\sim p_{\text{data}}(x,y)$, we specify the focus label $y_{\text{focus}}=y_{\text{focus}}(x,y,I_{\text{focus}})\in\mathcal{Y}$ that depends on the ground truth label $y$ and focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$. Intuitively, we define focus label $y_{\text{focus}}$ as $y_{\text{focus}}=y$ when either no focus features are specified (i.e., using the empty focus instruction), when the focus is on the underlying core feature $C$, or when ignoring a spurious feature $S$; but when either the focus is on a spurious feature or the core feature is ignored, $y_{\text{focus}}$ is defined as the label spuriously correlated with a particular value of the spurious feature present in input $x$. This changing target $y_{\text{focus}}$ trains the model to learn to adjust its responses based on specified features. Formally, we define $y_{\text{focus}}$ as:
$y_{\text{focus}}=\begin{cases}y&\text{if }I_{\text{focus}}\in\{\emptyset,\;\text{focus}(C),\;\text{focus}(C)\wedge\text{ignore}(S),\text{ignore}(S)\mid S\in\mathcal{S}\};\\ y_{s}&\text{if }I_{\text{focus}}\in\{\text{focus}(S),\;\text{focus}(S)\wedge\text{ignore}(F_{j})\mid F_{j}\in\mathcal{F}\setminus\{S\}\},\text{ for }s\in x;\\ y_{s}&\text{if }I_{\text{focus}}\in\{\text{ignore}(C)\}\text{ for sampled feature }S\in\mathcal{S}\text{ with value }s\in x.\\ \end{cases}$ (2)
where we again use $s\in x$ to denote the presence of the feature value $s$ in input $x$. See Figure 2 for a concrete example showing the focus label values for an example from the MNLI dataset under different focus instructions.
The objective of FIT training is to minimise the negative log-likelihood (NLL) of the response $y_{\text{focus}}$ conditioned on $I,I_{\text{focus}},x$. Formally, we define the FT loss objective as:
$\min_{\theta}\mathbb{E}_{(x,y)\sim p_{\text{data}}(x,y),\;I_{\text{focus}}\sim p_{\mathcal{I}_{\text{focus}}}(\mathcal{I}_{\text{focus}})}\;\left[-\log p_{\theta}\left(y_{\text{focus}}\mid I,I_{\text{focus}},x\right)\right].$ (3)
We define $p_{\mathcal{I}_{\text{focus}}}(\mathcal{I}_{\text{focus}})$ by placing a small probability mass on the empty focus instruction prompt $\emptyset$ in order to aid in learning the underlying task, and then uniformly distribute the remaining probability mass over the remaining non-empty feature instructions. The objective in Equation 3 can be optimised through sampling using stochastic gradient descent (SGD) with popular optimisers such as AdamW (Loshchilov & Hutter, 2019). Further details on FT optimisation are provided in Section A.1.
@@
Definition 1.
(Defining $\rho_{\text{spurious}}$) . Let $S\in\mathcal{S}\subseteq\mathcal{F}$ denote a spurious feature. Suppose that a value of $S$, say $s$, is spuriously correlated with label $y_{s}$. Then we define
$\rho_{\text{spurious}}(s)=\mathbb{P}(Y=y_{s}|S=s),$ (4)
where the $Y$ denotes a ground truth dataset label.
By varying $\rho_{\text{spurious}}(s)$, we can control the predictivity of spurious features and observe the model’s behaviour when focusing on or ignoring these features as well, as core features.
To enable the model to effectively learn to utilise focus instructions and predict focus labels, thereby inducing steerable behaviour, we require that the spurious feature $S$ is not predictive of the ground truth label $Y$ during training, that $Y\perp\!\!\!\perp S$. Similarly, we require that the core feature $C$ is not predictive of the spurious label $Y_{S}$, that is $Y_{S}\perp\!\!\!\perp C$. To enforce these independence conditions within our synthetic datasets, described in Section 4, we set $\rho_{\text{spurious}}=1/N$ for tasks with $N$ class labels and ensure a balanced label distribution during training. We prove that these conditions are sufficient to enforce $Y\perp\!\!\!\perp S$ in our generated synthetic SS and SMNLI training datasets, which are further detailed in Section A.9 and Section A.11. A more comprehensive discussion of this approach, including the effects of varying $\rho_{\text{spurious}}$ during training, is provided in Section A.1.
Next, we evaluate FIT across several test sets that capture different conditions of spurious correlations and distribution shifts:
@@
We further evaluate FIT under distribution shifts, where the specific values taken by spurious features do not overlap between the training and test sets, by introducing additional test sets:
[…]
We evaluate over these datasets specifically on our SMNLI dataset (c.f. Section 4.2).

### v3

@@ 1 Introduction
Instruction Tuning (IT) (Zhang et al., 2023), a specialised form of supervised fine-tuning (SFT), has become an essential step in the process of developing effective instruction-following large language models (LLMs) (Ouyang et al., 2022; Touvron et al., 2023; Chen et al., 2024). While extensive pre-training to perform next-token prediction allows LLMs to extract common patterns and knowledge from large text corpora, IT fine-tunes these models on input-output pairs complemented by natural-language task instructions, teaching them to perform open-ended language-based tasks given instructions (Huang et al., 2023).
@@
However, despite the improvements observed in zero-shot generalisation from IT, recent studies suggest that some of these gains may be superficial, stemming from the models’ ability to learn task template formats or spurious input/output correlations rather than a more generalisable instruction-following capability (Kung & Peng, 2023; Ghosh et al., 2024). As a result, LLMs may fail to generalise to new contexts where the same templates or spurious correlations are not present (Kung & Peng, 2023). Furthermore, fine-tuning can inadvertently lead to safety misalignment, where models lose alignment with desired objectives and become more prone to generating harmful or undesirable outputs (Qi et al., 2023). This motivates the need for methods that can adaptively steer models at inference time, enabling users to control model behaviours based on evolving requirements or safety considerations.
Figure 1: Focus Instruction Tuning (FIT). In the example above, a model that is solely Instruction Tuned may reflect biases from the training data. For instance, in a question from BBQ (Parrish et al., 2022), when asked who posed a technical question at an engineering candidate’s interview involving both a man and a woman, the model might incorrectly answer “the man” due to biases, despite evidence to the contrary. In contrast, a FIT model can ignore the gender feature and focus on the interview content, demonstrating steerability and adaptability at inference time.
@@
Accordingly, we propose Focus Instruction Tuning (FIT), an extension of traditional IT that fine-tunes LLMs with respect to an instruction specifying which features to “focus on” or “ignore.” FIT trains LLMs to condition responses based on these focus specifications and respond differently to the same task input based on the specified features, allowing end users to dynamically steer model behaviour. This capability provides precise, explainable control over features leveraged by models, and can be used to enforce desired invariances. For instance, in Figure 1, we illustrate how FIT can be used to steer a model to ignore gender stereotypes and focus on task-relevant information, enabling it to correctly solve a question-answering task.
@@
In our experiments, we demonstrate that FIT enables precise control over model behaviour by steering it to focus on task-relevant features while ignoring features specified as irrelevant or spurious (see Section 4). This flexibility allows FIT to address challenges such as mitigating the influence of demographic features in question-answering tasks, while also proving effective across diverse NLP tasks like sentiment analysis and natural language inference. Furthermore, we show that FIT is robust to distribution shifts over feature values and can generalise to new, unseen features, demonstrating its adaptability to dynamic user requirements and varying task contexts.
In summary, our primary contributions are as follows:
@@
    We introduce Focus Instruction Tuning (FIT), a method that allows users to flexibly and dynamically specify what features a model should or should not focus on when performing a task at inference time. FIT enables practitioners to incorporate domain-specific knowledge about core, spurious, or bias-relevant features in order to steer models according to the desired feature specification.
@@
    We show that focus tuning generalizes both to new features not seen during training, and to distribution shift over feature values.
@@ 2.1 Spurious Feature Learning
Deep neural networks, such as foundation models like LLMs, are susceptible to relying on spurious features present in the training dataset – i.e., input features that are correlated with outputs in the training distribution, but are not correlated in all test distributions (Ye et al., 2024). Relying on spurious features leads models to fail to generalise under distribution shifts where such correlations may no longer hold (Wang et al., 2023a). Spurious features have been extensively studied in computer vision, encompassing features such as background colour (Arjovsky et al., 2019; Xiao et al., 2021; Venkataramani et al., 2024; Hemmat et al., 2024), texture (Geirhos et al., 2018; Baker et al., 2018), or scene elements (Hemmat et al., 2024), and are also prevalent in many widely used NLP benchmarks (Sun et al., 2024; Borkan et al., 2019). For instance, the token SPIELBERG is spuriously correlated with positive sentiment in datasets like SST-2 (Socher et al., 2013b), meaning that models trained on SST-2 may learn to predict sentiment by leveraging these spurious features instead of more general sentiment features (Wang & Culotta, 2020). This reliance on non-causal features undermines the robustness of models in generalising to distribution shift. Traditional approaches for detecting and mitigating spurious feature learning, particularly under distribution shifts, include prompt engineering (Sun et al., 2024), regularisation techniques (Arjovsky et al., 2019; Chew et al., 2024), counterfactual inference (Wang & Culotta, 2020, 2021; Udomcharoenchaikit et al., 2022), or generating synthetic interventional data (Bansal & Grover, 2023; Yuan et al., 2024; Wang et al., 2024).
Mechanistic Interpretability.
Substantial work in mechanistic interpretability has also aimed to discover models’ latent representation of, and reliance on, various features (Davies & Khakzar, 2024). For instance, causal probing trains supervised probing classifiers to predict and modify feature representations encoded by foundation models Belinkov, 2022, and has been deployed to study how LLMs leverage task-causal versus spurious features (Ravfogel et al., 2021; Lasri et al., 2022; Davies et al., 2023; Canby et al., 2024). Other works have leveraged unsupervised mechanistic interpretability methods, such as circuit discovery techniques (Wang et al., 2023b; Conmy et al., 2023) and sparse auto-encoders (Subramanian et al., 2018; Yun et al., 2021), to improve generalisation by discovering spurious features leveraged by models in performing a given task and ablating their use of these features (Gandelsman et al., 2024; Marks et al., 2024). Finally, concept removal methods locate and manipulate supervised feature representations corresponding to bias features encoded by foundation models in order to remove these features (Ravfogel et al., 2020, 2022, 2023; Iskander et al., 2023; Belrose et al., 2024; Kuzmin et al., 2024).
@@ 2.2 Controlling LLMs
Aligning LLMs. Alignment techniques like Reinforcement Learning with Human Feedback (RLHF) (Bai et al., 2022) are powerful tools for aligning LLMs with annotated preference data and lead to reduced prevalence of harmful behaviour (Ouyang et al., 2022; Bai et al., 2022; Touvron et al., 2023; Korbak et al., 2023). However, RLHF-trained models still exhibit key alignment limitations such as sycophancy (defaulting to agreement with users even when incorrect or harmful; Perez et al., 2023; Sharma et al., 2024), and can still be adversarially prompted to generate harmful responses (Carlini et al., 2024). Furthermore, even well-aligned models can rapidly fall out of alignment when they are fine-tuned (Zhan et al., 2024; Yang et al., 2024; Lermen & Rogers-Smith, 2024), even on benign tasks (Qi et al., 2023). Thus, effectively steering model behaviours across the range of potential safety concerns that might emerge during LLM pre-training, fine-tuning, or deployment to novel contexts is an important and challenging goal in AI safety, necessitating more flexible and generalisable steering methods (Anwar et al., 2024).
Latent Steering.
A growing literature has worked to address this challenge via inference-time steering, where LLMs do not need to be retrained with respect to safety limitations, but can instead be controlled at inference time to steer models towards desirable behaviours or away from undesirable ones. For instance, latent steering methods perform embedding-space interventions that push models towards desirable behaviours or away from undesirable ones (Turner et al., 2023; Zou et al., 2023; Bhattacharjee et al., 2024; Li et al., 2024; Han et al., 2024). However, these methods require white-box access to model representations at inference time; and interventions are specific to each target behaviour, meaning that they must be re-computed in the context of any given target behaviour. Our work addresses these limitations by training models to control and adapt their responses based on explicit focus specifications, enabling users to flexibly and dynamically steer model behaviour at inference-time using simple, natural-language instructions.
@@ 3 Methodology
Preliminaries. We consider a pre-trained, decoder-only large language model (LLM) $p_{\theta}$ that models the probability of token sequences autoregressively over its vocabulary $\mathcal{V}$. Given a sequence of tokens $s=[s_{1},\dots,s_{L}]\in\mathcal{V}^{L}$, the jonit probability of $s$ under the model is given as
$p_{\theta}(s)\;=\;\prod_{i=1}^{L}\,p_{\theta}(s_{i}\mid s_{<i}),$ (1)
where $p_{\theta}(s_{1}\mid\emptyset)=p_{\theta}(s_{1})$. In supervised fine-tuning, we minimise the negative log-likelihood (NLL) of the output tokens $y$ given the input tokens $x$ using the autoregressive formulation defined in Equation 1.
In instruction tuning (IT) (Zhang et al., 2023), a form of SFT, an additional task instruction $I$ accompanies the input-output pair $(x,y)$, forming a tuple $(I,x,y)$. The objective becomes minimising the NLL of $y$ given both $I$ and $x$ over the distribution of input-output pairs, and instructions.
@@
[…]
$,\;{\color[rgb]{0,0.62109375,0.44921875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.62109375,0.44921875}\text{focus}(F_{i})\wedge\text{ignore}(F_{j})}\mid F_{i},F_{j}\in\mathcal{F}\},$
where: $\emptyset$ denotes an empty focus instruction with no features to focus on or to ignore; $\text{focus}(F_{i})$ is an instruction to focus on feature $F_{i}$; $\text{ignore}(F_{j})$ is an instruction to ignore feature $F_{j}$; and $\text{focus}(F_{i})\wedge\text{ignore}(F_{j})$ is an instruction to focus on feature $F_{i}$ whilst ignoring feature $F_{j}$. We include the default prompt in order to aid the model in learning the underlying task as well as the ability to refocus its attention on specified features during FIT. For specific examples of focus instructions that we consider, see Appendix E.
@@
Consider a sample $(x,y)\sim p_{\text{data}}$ drawn from an underlying data distribution, and a focus instruction $I_{\text{focus}}$ drawn from a distribution $p_{\mathcal{I}_{\text{focus}}}$ over the set of focus instructions $\mathcal{I}_{\text{focus}}$. Then the likelihood of response $y$ conditioned on input $x$, task instruction $I$ (as in standard IT), and focus-instruction $I_{\text{focus}}$ is modelled as $p_{\theta}(y|I,I_{\text{focus}},x)$.
@@
For a sample $(x,y)\sim p_{\text{data}}$, we specify the focus label $y_{\text{focus}}=y_{\text{focus}}(I_{\text{focus}},s,y)\in\mathcal{Y}$ that depends on the ground truth label $y$, focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$ and spurious feature value $s\in\text{Val}(S)$ present in $x$. Intuitively, we define focus label $y_{\text{focus}}$ as $y_{\text{focus}}=y$ when either no focus features are specified (i.e., using the empty focus instruction), when the focus is on the underlying core feature $C$, or when ignoring a spurious feature $S$; but when either the focus is on a spurious feature or the core feature is ignored, $y_{\text{focus}}$ is defined as $y_{\text{focus}}=y_{s}$, where $y_{s}$ is the label spuriously correlated with the particular spurious feature value $s$ present in input $x$. This changing target $y_{\text{focus}}$ trains the model to learn to adjust its responses based on feature specifications implemented through focus instructions. More formally, we define $y_{\text{focus}}$ as:
$y_{\text{focus}}=\begin{cases}y&\text{if }I_{\text{focus}}\in\mathcal{I}_{\text{focus}}^{C};\\ y_{s}&\text{if }I_{\text{focus}}\in\mathcal{I}_{\text{focus}}^{s}\;,\end{cases}$ (3)
where we define the task-casual and spurious instruction target sets as
$\mathcal{I}_{\text{focus}}^{y}$ =$=\{\emptyset,\;\text{focus}(C),\;\text{focus}(C)\wedge\text{ignore}(S),\;$ (4)
$\qquad\qquad\qquad\qquad\;\;\text{ignore}(S)\mid S\in\mathcal{S}\},$ (5)
$\mathcal{I}_{\text{focus}}^{y_{s}}$ =$=\{\text{focus}(S),\;\text{focus}(S)\wedge\text{ignore}(F_{j}),$ (6)
$\ \qquad\qquad\qquad\qquad\mid F_{j}\in\mathcal{F}\setminus\{S\}\},$ (7)
respectively. See Figure 2 for a concrete example showing the focus label values for an example from the MNLI dataset under different focus instructions.
The objective of FIT training is to minimise the negative log-likelihood (NLL) of the response $y_{\text{focus}}$ conditioned on $I,I_{\text{focus}},x$. Formally, as a form of expected-risk minimisation (ERM) (Vapnik et al., 1998), we define the FT loss objective as:
$\min_{\theta}\mathbb{E}_{(x,y)\sim p_{\text{data}},\;I_{\text{focus}}\sim p_{\mathcal{I}_{\text{focus}}}}\;\left[-\log p_{\theta}\left(y_{\text{focus}}\mid I,I_{\text{focus}},x\right)\right].$ (8)
We define $p_{\mathcal{I}_{\text{focus}}}(\mathcal{I}_{\text{focus}})$ by placing a small probability mass on the empty focus instruction prompt $\emptyset$ in order to aid in learning the underlying task, and then uniformly distribute the remaining probability mass over the remaining non-empty focus instructions. The objective in Equation 8 can be optimised through sampling using stochastic gradient descent (SGD) with popular optimisers such as AdamW (Loshchilov & Hutter, 2019). Further details on FT optimisation are provided in Appendix C.
@@
Definition 3.1.
(Defining $\rho_{\text{spurious}}$) . Let $S\in\mathcal{S}\subseteq\mathcal{F}$ denote a spurious feature. Suppose that a value of $S$, say $s\in\text{Val}(S)$, is spuriously correlated with label $y_{s}$. Then, for ground truth dataset label $Y$, we define
$\rho_{\text{spurious}}(s)=\mathbb{P}(Y=y_{s}|S=s)$ (9)
By varying $\rho_{\text{spurious}}(s)$, we can control the predictivity of spurious features and observe the model’s behaviour when focusing on or ignoring these features as well as core features under distribution shift.
To examine how well models can utilise focus instructions for steerable behaviour in a controlled environment, we construct synthetic datasets such that the spurious feature $S$ is not predictive of the ground-truth label $Y$ (i.e., $Y\perp\!\!\!\perp S$) and the core feature $C$ is not predictive of the spurious label $Y_{S}$ (i.e., $Y_{S}\perp\!\!\!\perp C$). We enforce these conditions by setting $\rho_{\text{spurious}}=1/N$ (for $N$ label classes) and ensuring a balanced label distribution during training. In Appendix I and Appendix K, we verify that our synthetic SS and SMNLI training sets introduced in Section 4.1 and Figure 4 respectively, satisfy these independence assumptions.
Next, we evaluate FIT across several test sets with varying predictivity levels:
@@
We further evaluate FIT under another form distribution shift, specifically on our SMNLI dataset (c.f. Section 4.2).. Here, the specific values taken by spurious features do not overlap between the training and test sets.
  • $\mathcal{D}^{s}_{\cdot}$: Test datasets where the spurious feature values are distinct from those within the training set. Here, we use the same predictivity levels as in the initial datasets presented above.
Note that, while we define FIT with respect to annotated spurious features, this requirement can be alleviated by, e.g., combining FIT with automated spurious feature identification methods (Wang et al., 2022; see Appendix B for further discussion).